This is a chapter over systems and matrices.. The section is over linear equations.
Here is the question
A store manager wishes to make a 100 pounds of nut mixture he can sell for $4.90 per pound. He must mix together peanuts at $3.00 per pound,cashews at $5.00 per pound , and pecans at $8.00 per pound. The pounds of peanuts must equal the amount of cashews and pecans combined. Find the amount of each type of nut required. Show all work, use algebra not matrices
Hope you can help!

Answer by richwmiller(17219)

This belongs in category :word problems:mixtures
n+c+p=100
3n+5c+8p=100*4.9
n=c+p
We now have three unknowns and three equations.
We should be able to solve this.
cashews= c=20, peanuts=n 50., pecans=p 30.